The column is composed of a W12x50 and two channel C6x13 in grade 50, length 16ft. Calculate rx, ry. resistance?

C6x 13 W12 x 50 16
Answer #1

W 12 X 50 C6613 t -x/d I In: 391 in Ty? 56.3 in Ag = 143 in y L 1 * & Inz 17.3 in Iqz 1.05 in A² 3.81 7 = 0.514in. bg= 2.16Since ruchy buckling about X-amis. goverry KL= 16ft (KL/P) = (16x12/2.036) = 94.3 Eulers Buckling stress Fe= TE/KL/4² = fe -